i need help with this problem
Stephen has $40,000 to invest in two accounts and one account pays 4% and the other pays 6%. The annual income from these accounts is $2040. How much did he invest in each account

Answer by richwmiller(17219)

Total amount of money invested: $40000
x+y=40000,
Total yearly interest for the two accounts is: $2040
0.04*x+0.06*y=2040
x=40000-y
Substitute for x
0.04*(40000-y)+0.06*y=2040
Multiply out
1600-0.04*y+0.06*y=2040
Combine like terms.
0.02*y=440
Isolate y
y=$ 22000.00 at 6% earns $1320 interest
x=40000-y
Calculate x
x=$ 18000.00 at 4% earns $720 interest
Check
0.04*18000+0.06*22000=2040
720+1320=2040
2040=2040
If 2040=2040 is TRUE and neither x nor y is negative then all is well
codeint1